OpenStudy (anonymous):

Solve for x: 4 / 3x + 2 / 3 = 6

OpenStudy (anonymous):

4 / 3x + 2 / 3 = 6 -2/3 -2/3 4/3x = 6-2/3 now?

OpenStudy (anonymous):

4 / 3x + 2 / 3 = 6 4 / 3x = 6- 2/3 x = (6- 2/3)/ (4/3x) how???

OpenStudy (anonymous):

(6- 2/3)/ (4/3x)?

OpenStudy (mayankdevnani):

4/3x=6*3/2=9 27x=4 x=27/4

OpenStudy (mayankdevnani):

ok @Stephhaniiee

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

4 / 3x + 2 / 3 = 6 -2/3 -2/3 4/3x = 6-2/3 4/3x = 16/3 x = 16/3*(3/4) x = 16/4 =4

OpenStudy (anonymous):

2/3(2/x +1)= 6 2/x+1 = 18/2 2/x=8 x=1/4